OpenStudy (anonymous):

Suppose x coins are tossed. Write an expression to represent the number of possible outcomes. A. 2^X B. X+X C. X*X D. 2X

OpenStudy (andriod09):

so waht do you know about this already?

OpenStudy (anonymous):

There are 2 sides to a coin

OpenStudy (andriod09):

>.< besides that. i mean what do you know about probability?

OpenStudy (anonymous):

I know the probability of a coin landing on heads is 1/2 the same for tails

OpenStudy (anonymous):

and I know every time you add a coin the chance that there will only be one heads halves

OpenStudy (andriod09):

so then you just answered your own question. what's the answer then?

OpenStudy (anonymous):

but its talking about outcomes

OpenStudy (anonymous):

not probability

OpenStudy (andriod09):

-_- >.< their the same region of math. >_<

OpenStudy (anonymous):

They might be the same region but even in the same region you can get very different answers like when using sin cos and tan

OpenStudy (anonymous):

Im confused here

OpenStudy (andriod09):

\[x^2\] that's the answer

OpenStudy (anonymous):

thats not one of the answers available

OpenStudy (kropot72):

 The diagram shows the outcomes of tossing 1 coin, 2 coins and 3 coins. There are 2 outcomes for 1 coin, 4 outcomes for 2 coins and 8 outcomes for 3 coins. Can you make the correct choice of expression now?

OpenStudy (anonymous):

um maybe

OpenStudy (kropot72):

Try substituting in each of the choices for 3 coins and see which gives the correct number of possible outcomes. For example choice B: x + x = 3 + 3 = 6 Therefore choice B is incorrect.

OpenStudy (kropot72):

@bookaholic Have you tried another substitution?

OpenStudy (kropot72):

@bookaholic Are you there?